OpenStudy (anonymous):

if 10% of a herd of 10,000 deer will die out durning the winter,how many deer will still remain?

OpenStudy (anonymous):

i think it is 1,000 but i want to be sure

OpenStudy (lucaz):

the remaining must be 100%-10%=90% right? 90% of 10,000 = 90/100 * 10,000

OpenStudy (anonymous):

so i was right

OpenStudy (anonymous):

90% of 10,000 = .9(10,000) = you do the math

OpenStudy (lucaz):

90/100*10,000 = 0.9*10,000 = 9,000, you calculated how many died, the 10%, if you subtract that from the 10,000 you get what the problem is asking

OpenStudy (anonymous):

its asking how many remain...so 1000 is not the answer

OpenStudy (anonymous):

9,000 because you times 9 times 10,000

OpenStudy (anonymous):

correct...9000 remain

OpenStudy (anonymous):

you multiply 0.9 x 10000